OpenStudy (topofmorning89):

For every box of Oreo taken two replace it. If there was 48 boxes of Oreos taken how many Oreos are there now?

OpenStudy (topofmorning89):

Do I just 48*2?

OpenStudy (tkhunny):

Insufficient information. We have 2 * 48 = 96 boxes added, but we do not know the number of boxes at the start. If we start with just 1 box, we get: 1 - 1 + 2 = 2 2 - 1 + 2 = 3 etc. 47 - 1 + 2 = 48 48 - 1 + 2 = 49 We MUST know the start number.